a man goes 10 m towards north and 20 m towards east. find the displacement.
A   22.5 m
B   25 m
C   25.5 m
D   30 m

Grade:11

3 Answers

Aman gohel
6 years ago
(10m)^------------------->(20m)||||According to triangle`s law of vector addition,=>R=√(10^2+20^2)=>R=√500=>R=22.5m(Ans)Therefore correct option is A.
sachin kumar
6 years ago
If we take east as x-axis and north as y-axis,then displacement =20i^+10j^. So , magnitudeof displacement a=√400+100=10√5=22.5m .
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
If we take east as x-axis and north as y-axis,then displacement
= 20i + 10j.
So, magnitude of displacement a = √(400+100) = 10√5 = 22.5m
 
Thanks and regards,
Kushagra